OpenStudy (anonymous):

Algebra 2 Test Help? Someone please help me ASAP! :(

OpenStudy (anonymous):

I'm not trying to cheat! I'm just dumb and completely failing Algebra.. Someone please take the time to help me :/ Which set of ordered pairs in the form of (x, y) does not represent a function of x? {(–1, 2), (3, –2), (0, 1), (5, 2)} {(–1, 2), (3, 2), (–2, 2), (0, 2)} {(–1, 2), (3, –2), (0, 1), (3, 5)} {(–1, 2), (2, 3), (3, –2), (–2, 0)}

OpenStudy (johnweldon1993):

Alright...just remember a function is only a function if it passes the vertical line test. This means for each value 'x' you can only have 1 unique 'y' value, you can never have 2 different 'y' for one x.....say you have something like (1,2) (1,3) This would look like  If I told you to tell me what value of 'y' is when x = 1 ...would you tel me 2 or 3? you wouldn't know..this is why it is not a function.. So look at all your possible choices...if you see an 'x' that appears more than once in the function...it fails, andin this case is your answer

OpenStudy (anonymous):

@johnweldon1993 thank you! can you help me with a few more hard ones?

OpenStudy (johnweldon1993):

Sure :)

OpenStudy (anonymous):

Which of the following equations best represents the regression line for the data given in the table? y = x + 2 y = 2x – 2 y = –x – 2 y = x – 2 x|y 1|0 2|-1 4|0 6|4 8|8 10|9 11|8

OpenStudy (johnweldon1993):

Okay this one we can play with...see how you have points? (1 , 0) (2 , -1) (4 , 0) etc... Well you have choices of different equations....so lets try one out...look at equation 1 y = x + 2 Take one of your points...(1 , 0) and plug it in 0 = 1 + 2 Is that correct? of course not... and just try on like that from there

OpenStudy (johnweldon1993):

Now you'll see that none of the equations actually solve this function...makes sense....but which one is the closest to solving it?

OpenStudy (anonymous):

is it the second choice?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Ehh...look at option 4 again ....looks alot closer to me

OpenStudy (anonymous):

but then if you plug in the first one (0,1) the equation is 0=1-2 and that's false? I'm confused:(

OpenStudy (johnweldon1993):

Right...read a couple posts up...the equation wont be solved....it will be close... Remember your question asks for a line of regression (line of best fit without actually being the solution line)

OpenStudy (johnweldon1993):

so what you want to do is find which equation makes the most of the points be true

OpenStudy (anonymous):

oh i see.. so it is answer 4?

OpenStudy (anonymous):

And here's my next question.. only one more after this and I'm finished :) Which equation defines the graph of y = x² after it is shifted vertically 3 units down and horizontally 5 units left? y = (x – 5)² – 3 y = (x + 5)² – 3 y = (x – 3)² – 5 y = (x + 3)² – 5

OpenStudy (anonymous):

oh no don't disappear now:(

OpenStudy (johnweldon1993):

lol no just helping someone else out too :)

OpenStudy (johnweldon1993):

Okay to shift a graph up or down...you add or subtract a number from the whole function to shift a graph left or right...you add or subtract a number from the 'x' So...we start with x^2 To shift this 3 units down...we subtract 3 from the function....so y = x^2 - 3 make sense? And to shift the graph to the left 5 units...you ADD 5 to the 'x' y = (x + 5)^2 - 3 Make sense?

OpenStudy (anonymous):

Yep! :) Here's the last one i needed help with... What is the graph of the absolute value equation? y = 5|x|

OpenStudy (johnweldon1993):

Okay this is the last transformation...a shrink or stretch situation...when you multiply a function by a number greater than 1...it shrinks it horizontally...when you multiply by a number less than 1...you guessed it...it stretches it horizontally...so where are you answer choices?

OpenStudy (anonymous):

Well, it was shaped like a v, but I started trying to solve that one and figured it out lol:) I turned it in and gor a 15/15 on my test!:)

OpenStudy (johnweldon1993):

So then you're perfect! :) great job!

OpenStudy (anonymous):

Thanks for the help!:)

OpenStudy (johnweldon1993):

Anytime :)